# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v2

@@ II-A2 Distributed Devices
Some studies utilize additional WiFi devices to enhance sensing range and improve generalization to human orientation and position. For instance, ReWiS [66], Widar 3.0 [67], and OneFi [68] adopt configurations of 1 transmitter with 6 receivers and 1 transmitter with 4 receivers, respectively, to extract environment-independent features. WiTraj [69] leverages 3 receivers placed at different viewing angles to capture human walking and achieve robust trajectory reconstruction. Additionaly, Wang et al. [245] propose a theoretical model analyzing how the distance between the transmitter and receiver influences the sensing range. Optimizing the placement of these devices effectively enhances the sensing area and reduces the impact of environmental interference.
@@ II-A3 Scaling up dataset
Scaling up training datasets is another effective approach to enhancing generalization capability, as demonstrated in Widar 3.0 [67], Zhang et al. [71], MM-Fi [72], XRF55 [70] and CSI-Bench [73]. For example, XRF55 [70] shows that training on approximately 30,000 samples enables the sensing model to naturally achieve direction and position invariance, while also supporting adaptation to new environments through few-shot fine-tuning.
@@ II-B1 Signal Analysis Methods
Signal analysis plays a crucial role in Wi-Fi sensing by extracting relevant features from raw signals, ensuring generalizability across various environments, locations, and orientations. WiHand [74] employs a Low Rank and Sparse Decomposition (LRSD) algorithm to separate gesture signals from background noise, enhancing its resilience to location variations. CLAR [75] utilizes Singular Spectrum Analysis (SSA) [246] to extract the trend components from Wi-Fi signals, enabling robust activity recognition. Gait-Enhance [76] introduces a CSI reconstruction algorithm that mitigates the negative impact on recognition accuracy caused by varying CSI patterns associated with walking in different directions. Zhuo et al.  [77] apply a combination of Principal Component Analysis (PCA) and Variational Mode Decomposition (VMD) [247] to reduce noise interference in non-ideal sleep positions, enabling position-free breath detection. Abuhoureyah et al. [78] leverage Independent Component Analysis (ICA) and Continuous Wavelet Transform (CWT) to decompose CSI signals, improving multi-user human action recognition in complex environments.
@@ II-B3 Motion Indicators
Many works compute domain-independent motion indicators in speed, velocity, direction, and trajectory, to enhance the generalizability of Wi-Fi sensing systems. Sharp [91], OneSense [92], Yin et al. [93], Wi-learner [94], Niu et al. [95], and LAGER [96] focus on utilizing Doppler Frequency Shift (DFS). Mini-Batch Alignment [98, 99] incorporates both DFS and Gramian Angular Difference Field (GADF). WiLife [97] further explores Doppler velocity as a key motion indicator. WiTraj [69] identifies speed ambiguity in the Doppler-MUSIC algorithm’s spectrum. To address this, it employs multiple receivers at different viewing angles and leverages the CSI quotient to eliminate this ambiguity, enabling robust trajectory reconstruction.
@@ II-B4 Angle Indicators
Angle indicators are also used to enhance the generalizability of Wi-Fi sensing. Li et al. [114] use Angle Difference of Arrival (ADoA) for location-free CSI-based activity recognition. EasyTrack [115] utilizes AoA–AoD values, leveraging changes in the length and direction of the Wi-Fi signal path. Through incremental calculations and dynamic path extraction, it can reconstruct gesture trajectories in real-time without relying on the initial gesture position or a fixed device location. 3D-ID [116] uses 2D AoA estimation of signal reflections to enable Wi-Fi to visualize a person in the environment. This visualization is then digitized into a 3D body representation using deep learning, which extracts both the static body shape and dynamic walking patterns for person re-identification.
@@ II-C1 Domain Alignment
Figure 7: Domain alignment with domain discriminator or domain classifier.
@@
Several variants have been proposed to enhance this domain-adversarial learning framework further. For example, i-Sample [120] adopts a two-stage training strategy. In the first stage, the model focuses on learning a domain-invariant feature extractor by jointly feeding source and target domain data. A domain discriminator is used to classify the domain origin of each sample, while a parallel branch performs activity classification. In the second stage, the feature extractor is frozen, and the classifier is fine-tuned using labeled source data. This staged approach effectively enhances cross-environment generalization performance. In another line of work, JADA [121] employs two separate encoders: one for the source domain and another for the target domain. Through adversarial training with a domain discriminator, the features extracted from both encoders are encouraged to become domain-invariant. After training, the source-domain encoder is frozen, and a classification head is trained using labeled source data. The final model—comprising the target-domain encoder and the trained classification head—is used for inference on target-domain samples, achieving robust cross-domain activity recognition. Other examples are CrossGR [122] and Wi-Adaptor [123], which replace the traditional domain discriminator with an auxiliary action classifier. Unlike standard classifiers, this branch is trained to assign uniform probabilities across all action classes, effectively making actions indistinguishable in the feature space. This strategy forces the feature extractor to remove task-relevant cues from the domain-variant features, thereby encouraging it to learn more domain-invariant and task-discriminative representations. As a result, the main action classifier benefits from cleaner, more generalizable features during training.
@@
In addition to MMD, several alternative metrics have been adopted to quantify distributional similarity between source and target domains. For example, Wi-CHARS [142] applies Earth Mover’s Distance (EMD) to align features for cross-environment activity recognition. Mehryar [143] utilizes the Kullback–Leibler (KL) divergence to reduce domain shift, while AdapLoc [144] employs Euclidean distance for indoor localization across domains. PIAS [145] further explores Contrastive Domain Discrepancy (CDD) [146] to align representations in a contrastive manner. WiCAU [147] uses Wasserstein distance to measure to quantify distributional similarity between source and target domains.
@@
Figure 9: Generative Adversarial Networks
Another line of research explores Generative Adversarial Networks (GANs) [252] to perform domain alignment by transforming data distributions or learning shared latent representations. GAN-based alignment builds on the adversarial learning mechanism of GANs, shown in Fig. 9, which involve a generator and a discriminator engaged in a minimax game. The generator aims to produce data indistinguishable from real samples, while the discriminator attempts to differentiate between real and generated data. Through iterative optimization, the generator learns to synthesize data that closely follows the expected distribution. For example, MetaGanFi [154] proposes a three-stage pipeline: (1) In the first stage, a CycleGAN-based framework [253] is used to convert Wi-Fi data from multiple source domains into a unified Uni-domain representation. This is achieved through adversarial training, where the generator transforms source-domain data to resemble a common distribution, and a discriminator is trained to identify the originating domain. The generator is optimized to fool the discriminator, thereby removing domain-specific cues. (2) In the second stage, the generated Uni-domain data are used to train an activity classifier, ensuring that the classifier learns from domain-invariant representations. (3) Finally, during inference, target-domain data are passed through the trained generator and classifier to perform activity recognition for unseen users. Similarly, Zhang et al. [155] and Wi-Cro [156] and WiTeacher [157] employ CycleGAN and StyleGAN [254], respectively, but instead of mapping source domains to a unified representation, they transform source domain data to match the target domain distribution, thereby enabling cross-domain sign language recognition and activity recognition.
@@ II-C3 Metric Learning for Zero/Few-shot Learning
Figure 11: Siamese Neural Networks
Siamese neural networks [263] offer a classical yet effective architecture for few-shot learning and domain generalization. In this framework, as shown in Fig. 11, a pair of samples is fed into two identical networks that share weights. Each network extracts feature embeddings from its input, and the resulting embeddings are concatenated and passed through a binary classifier. The classifier is trained to predict whether the two inputs belong to the same class: if they do (e.g., two samples of the same activity), the target label is 1; otherwise, it is 0. During inference, a support set is constructed from a few labeled samples per class in the target domain. To classify a query sample, it is paired with each support sample and processed by the Siamese network. The network outputs a confidence score indicating the likelihood of the two samples belonging to the same class. The class label of the support sample with the highest confidence score is then assigned to the query sample.
This pipeline has been widely adopted in Wi-Fi sensing tasks for few-shot cross-domain generalization. For instance, 3D-ID [116] utilizes this pipeline for few-shot user identification in unseen environments. MaP-SGAN [178] applies it for gait recognition by enhancing the diversity of training samples via generative models. CrossFi [179] extends the Siamese framework to handle both cross-domain and new-class scenarios, enabling gesture recognition and user identification in previously unseen domain.
@@
Figure 12: Relation Network
@@
Figure 13: Matching Network
@@
Figure 14: Prototypical Network
@@
Figure 15: Model-Agnostic Meta-Learning (MAML) [194] includes a meta-training stages and a meta-testing stage. The meta-training stage is designed to find a good initialization of model parameters such that the model can quickly adapt to new tasks with only a few gradient steps and a small amount of labeled data in the meta-testing stage.

### v4

@@ II-A Experimental Setup
To explicitly connect the challenges identified in Fig. 2, we note that each solution is designed to mitigate specific forms of diversity in Wi-Fi sensing systems.
  • Device heterogeneity, arising from variations in hardware capabilities, antenna configurations, and radio characteristics across commercial Wi-Fi devices, is primarily addressed through distributed antennas and distributed devices setups, which exploit spatial and hardware diversity to improve robustness.
  • Human body diversity, including differences in body shape, orientation, motion patterns, and carrying conditions, is mitigated by distributed antennas, as multi-view signal observations reduce sensitivity to user orientation and occlusion, and by scaling up datasets, which enables models to generalize across diverse human characteristics.
  • Environment diversity, caused by changes in room layout, materials, multipath profiles, and deployment locations, is addressed by distributed devices and optimizing device placement that capture spatially heterogeneous channel responses, as well as large-scale and multi-environment datasets, which expose learning models to a broad range of propagation conditions.
Together, these experimental setup strategies form a systematic approach to overcoming the key diversity challenges in real-world Wi-Fi sensing.
@@ II-A2 Distributed Devices
Some studies utilize additional Wi-Fi devices to enhance sensing range and improve generalization to human orientation and position. For instance, ReWiS [9], Widar 3.0 [313], and OneFi [244] adopt configurations of 1 transmitter with 6 receivers and 1 transmitter with 4 receivers, respectively, to extract environment-independent features. WiTraj [235] leverages 3 receivers placed at different viewing angles to capture human walking and achieve robust trajectory reconstruction. Additionally, Wang et al. [219] propose a theoretical model analyzing how the distance between the transmitter and receiver influences the sensing range. Optimizing the placement of these devices effectively enhances the sensing area and reduces the impact of environmental interference.
@@ II-A3 Scaling up dataset
Scaling up training datasets, such as collecting large-scale dataset and data augmentation, is another effective approach to enhancing generalization capability, as demonstrated in Widar 3.0 [313], Zhang et al. [304], MM-Fi [254], XRF55 [208], and CSI-Bench [323]. For example, XRF55 [208] shows that training on approximately 30,000 samples enables the sensing model to naturally achieve direction and position invariance, while also supporting adaptation to new environments through few-shot fine-tuning.
Table II compares the advantages and limitations of three experimental setup strategies, explains their underlying generalization mechanisms, and summarizes the scenarios in which each approach is best suited. Based on this comparison, the analysis of experimental setups provides several strategic insights into building generalizable Wi-Fi sensing systems.
Key Takeaways and Lessons Learned: Experimental Setup Structural Invariance through Spatial Diversity: Distributed antenna topologies (e.g., staggered or circular layouts) and multi-device configurations (e.g., 1-TX with 4-6 RXs) provide a physical-layer solution to orientation sensitivity. This “structural invariance” ensures multi-view signal observations that effectively mitigate body-shadowing and occlusion. Placement as a Robustness Multiplier: Optimizing the geometric placement of transmitters and receivers is critical for capturing spatially heterogeneous channel responses. Strategic layout design not only maximizes the sensing range but also significantly reduces the impact of environmental multipath interference. Natural Invariance through Data Scaling: A key lesson from large-scale datasets (e.g., XRF55, CSI-Bench) is the existence of a quantitative threshold where models begin to exhibit “natural” invariance. Massive data exposure allows models to inherently decouple human motion from specific propagation conditions, facilitating easier adaptation to novel domains. Exploiting Diversity to Counter Heterogeneity: To overcome device heterogeneity, moving from single-link to multi-link coordinated sensing is essential. Distributed setups exploit spatial and hardware diversity to average out the unique radio artifacts and hardware imperfections of individual commercial Wi-Fi chipsets.
@@
TABLE III: Comparison of Signal Preprocessing Techniques for Wi-Fi Sensing Generalization.
Category Pros (Advantages) Cons (Limitations) Generalization Mechanism Best-fit Scenarios
Signal Analysis ([141, 317, 255, 327, 2]) Effective for noise removal, trend extraction, and background subtraction. Often computationally intensive; may lose non-linear motion details. Extracting motion-related components via signal decomposition (e.g., PCA, ICA, SSA). Stable environments; offline analysis; scenarios requiring strong denoising.
Handcrafted ([149, 88, 57, 281, 263, 302, 157, 220, 47, 158, 20, 274]) High interpretability; robust to specific interference (e.g., CCI) and phase errors. Sensitive to hardware variations and chipset-specific phase noise. Designing invariant indicators (e.g., Phase Difference, IQ curvature) to bypass shifts. Resource-constrained systems; applications demanding interpretability and robustness.
Motion Indicators ([148, 306, 261, 52, 154, 25, 122, 194, 195, 235, 313, 272, 256, 181, 22, 299, 244, 32, 58, 179, 225, 236, 211, 165, 288, 289, 276, 324]) High-level physical abstraction; achieves environment and location independence. Requires high-resolution CSI or specific multi-antenna configurations. Mapping signals to body-coordinate velocity/speed profiles (e.g., BVP, DFS). Cross-environment deployment; user-independent sensing; mobility-centric tasks.
Angle Indicators ([127, 73, 172]) Provides precise spatial geometry; robust for trajectory tracking and Re-ID. Performance is sensitive to antenna array calibration and layout. Utilizing spatial consistency (e.g., AoA, ADoA) to digitize human posture and movement. Localization, tracking, and identification tasks with calibrated antenna arrays.
@@ II-B1 Signal Analysis Methods
Signal analysis plays a crucial role in Wi-Fi sensing by extracting relevant features from raw signals, ensuring generalizability across various environments, locations, and orientations. WiHand [141] employs a Low Rank and Sparse Decomposition (LRSD) algorithm to separate gesture signals from background noise, enhancing its resilience to location variations. CLAR [317] utilizes Singular Spectrum Analysis (SSA) [61] to extract the trend components from Wi-Fi signals, enabling robust activity recognition. Gait-Enhance [255] introduces a CSI reconstruction algorithm that mitigates the negative impact on recognition accuracy caused by varying CSI patterns associated with walking in different directions. Zhuo et al.  [327] apply a combination of PCA and Variational Mode Decomposition (VMD) [48] to reduce noise interference in non-ideal sleep positions, enabling position-free breath detection. Abuhoureyah et al. [2] leverage Independent Component Analysis (ICA) and Continuous Wavelet Transform (CWT) to decompose CSI signals, improving multi-user human action recognition in complex environments.
@@ II-B3 Motion Indicators
Many works compute domain-independent motion indicators in speed, velocity, direction, and trajectory, to enhance the generalizability of Wi-Fi sensing systems. Sharp [148], OneSense [306], Yin et al. [261], Wi-learner [52], Niu et al. [154], and LAGER [25] focus on utilizing Doppler Frequency Shift (DFS). Mini-Batch Alignment [194, 195] incorporates both DFS and Gramian Angular Difference Field (GADF). WiLife [122] explores Doppler velocity as a key motion indicator. WiTraj [235] identifies speed ambiguity in the Doppler-MUSIC algorithm’s spectrum. To address this, it employs multiple receivers at different viewing angles and leverages the CSI quotient to eliminate this ambiguity, enabling robust trajectory reconstruction.
@@ II-B4 Angle Indicators
Angle indicators are also used to enhance the Wi-Fi sensing generalizability. Li et al. [127] use Angular Difference of Arrival for location-free CSI-based activity recognition. EasyTrack [73] utilizes AoA–AoD values, leveraging changes in the length and direction of the Wi-Fi signal path. Through incremental calculations and dynamic path extraction, it can reconstruct gesture trajectories in real-time without relying on the initial gesture position or a fixed device location. 3D-ID [172] uses 2D AoA estimation of signal reflections to enable Wi-Fi to visualize a person in the environment. This visualization is then digitized into a 3D body representation, which extracts both the static body shape and dynamic walking patterns for person re-identification.
Table III compares the advantages and limitations of four signal preprocessing strategies, explains their underlying generalization mechanisms, and summarizes the scenarios in which each approach is best suited. Based on this comparison, the analysis of signal preprocessing provides several critical takeaways and lessons learned for building generalizable Wi-Fi sensing systems.
Key Takeaways and Lessons Learned: Signal Preprocessing Signal Analysis for Denoising: Algorithms like LRSD, SSA, and PCA are fundamental for extracting action-related components from background noise. These methods enhance resilience to location variations and enable robust monitoring (e.g., breath detection) even in non-ideal positions. Handcrafted IQ & Phase Features: Advanced indicators derived from phase differences or the IQ complex plane capture fine-grained patterns. Notably, features like NLPEV provide independence CCI, which is crucial for real-world robustness. Physical Invariance via Motion Indicators: Transitioning to speed/velocity profiles, particularly BVP and the Fresnel Zone Model, represents a shift toward domain invariance. By mapping signals to speed or body-relative coordinates, systems decouple sensing from user orientation and environment. Geometric Consistency with Angle Indicators: Angle-based measurements, such as AoA and ADoA, provide spatial geometric paths consistent across domains. These enable real-time trajectory reconstruction and 3D body representation, supporting complex tasks like person re-identification. Signal-to-Visual Feature Enrichment: Converting 1D CSI time-series into 2D representations (e.g., Recurrence Plots, Grayscale images) bridges the gap between signal processing and computer vision, allowing models to benefit from powerful pre-trained visual architectures and data augmentation.
@@ II-C1 Domain Alignment
Several variants have been proposed to enhance this domain-adversarial learning framework further. For example, i-Sample [321] adopts a two-stage training strategy. In the first stage, the model focuses on learning a domain-invariant feature extractor by jointly feeding source and target domain data. A domain discriminator is used to classify the domain origin of each sample, while a parallel branch performs activity classification. In the second stage, the feature extractor is frozen, and the classifier is fine-tuned using labeled source data. This staged approach effectively enhances cross-environment generalization performance. In another line of work, JADA [329] employs two separate encoders: one for the source domain and another for the target domain. Through adversarial training with a domain discriminator, the features extracted from both encoders are encouraged to become domain-invariant. After training, the source-domain encoder is frozen, and a classification head is trained using labeled source data. The final model, comprising the target-domain encoder and the trained classification head, is used for inference on target-domain samples, achieving robust cross-domain activity recognition. Other examples are CrossGR [125] and Wi-Adaptor [278], which replace the traditional domain discriminator with an auxiliary action classifier. Unlike standard classifiers, this branch is trained to assign uniform probabilities across all action classes, effectively making actions indistinguishable in the feature space. This strategy forces the feature extractor to remove task-relevant cues from the domain-variant features, thereby encouraging it to learn more domain-invariant and task-discriminative representations. As a result, the main action classifier benefits from cleaner, more generalizable features during training.
Figure 7: Domain alignment with domain discriminator or domain classifier. Both source and target data are mapped into a shared feature space. Through adversarial competition between an action classifier and a domain discriminator, the network learns to produce domain-agnostic representations that enhance cross-domain performance.
@@
In addition to MMD, several alternative metrics have been adopted to quantify distributional similarity between source and target domains. For example, Wi-CHARS [294] applies Earth Mover’s Distance (EMD) to align features for cross-environment activity recognition. Mehryar [146] utilizes the Kullback–Leibler (KL) divergence to reduce domain shift, while AdapLoc [318] employs Euclidean distance for indoor localization across domains. PIAS [245] further explores Contrastive Domain Discrepancy (CDD) [100] to align representations in a contrastive manner. WiCAU [34] uses Wasserstein distance to measure and quantify distributional similarity between source and target domains.
@@
Figure 9: Generative Adversarial Networks. Generator ($G$): The network responsible for creating synthetic data. Its primary objective is to learn the underlying distribution of the real data and produce samples that are indistinguishable from the ground truth to “fool” the discriminator. Discriminator ($D$): The network acting as a binary classifier. Its role is to evaluate both real data ($x$) and synthetic data ($G(z)$), outputting a probability that a given sample is “real” (from the actual dataset) rather than “fake” (produced by the generator). It provides the gradient signal that guides the generator to improve its realism.
Another line of research explores Generative Adversarial Networks (GANs) [63] to perform domain alignment by transforming data distributions or learning shared latent representations. GAN-based alignment builds on the adversarial learning mechanism of GANs, shown in Fig. 9, which involves a generator (G) and a discriminator (D) engaged in a minimax game. The generator aims to produce data indistinguishable from real samples, while the discriminator attempts to differentiate between real and generated data. Through iterative optimization, the generator learns to synthesize data that closely follows the expected distribution. For example, MetaGanFi [284] proposes a three-stage pipeline: (1) In the first stage, a CycleGAN-based framework [326] is used to convert Wi-Fi data from multiple source domains into a unified Uni-domain representation. This is achieved through adversarial training, where the generator transforms source-domain data to resemble a common distribution, and a discriminator is trained to identify the originating domain. The generator is optimized to fool the discriminator, thereby removing domain-specific cues. (2) In the second stage, the generated Uni-domain data are used to train an activity classifier, ensuring that the classifier learns from domain-invariant representations. (3) Finally, during inference, target-domain data are passed through the trained generator and classifier to perform activity recognition for unseen users. Similarly, Zhang et al. [279] and Wi-Cro [145] and WiTeacher [242] employ CycleGAN and StyleGAN [102], respectively, but instead of mapping source domains to a unified representation, they transform source domain data to match the target domain distribution, thereby enabling cross-domain sign language recognition and activity recognition.
@@
Summary and Design Insights for Domain Alignment
The above domain alignment strategies reveal a fundamental trade-off between alignment granularity, supervision availability, and training stability. Adversarial alignment with domain discriminators provides a flexible and widely adopted solution for unsupervised domain adaptation, particularly when only source-domain labels are available. Extending the discriminator to multi-class domain classifiers further enables disentanglement across multiple domain factors (e.g., user, environment, orientation), making it well-suited for complex real-world deployments with explicitly labeled domain attributes. However, adversarial training may introduce optimization instability and potential over-alignment, motivating the exploration of more stable alternatives.
Similarity-based approaches, such as MMD, Wasserstein distance, and KL divergence, offer a theoretically grounded and optimization-friendly mechanism for distribution matching. They are particularly effective under moderate domain shifts and limited computational budgets, though their effectiveness may diminish when distribution gaps become large. When finer-grained supervision is available, class-level or sub-domain alignment provides more precise cross-domain matching, reducing negative transfer by aligning semantically corresponding categories instead of global distributions. Nevertheless, such approaches rely on accurate class labels or high-quality pseudo-labels, which may not always be obtainable.
GAN-based alignment extends domain adaptation by explicitly transforming data distributions, making it advantageous for large domain discrepancies or cross-modality translation. However, its higher computational complexity and adversarial instability limit scalability in resource-constrained sensing systems. Multi-task learning alignment, in contrast, improves generalization by leveraging diversity across tasks or domains without necessarily requiring target-domain data, making it attractive for domain generalization settings. Finally, cross-modal embedding represents a promising direction, as it connects Wi-Fi sensing representations with semantically rich modalities (e.g., text or radar), enabling zero-shot generalization and integration with foundation models.
In practice, the selection of domain alignment strategy should be guided by (i) the availability of domain and class labels, (ii) the magnitude and nature of domain shifts, and (iii) computational constraints. It is expected to move toward hybrid frameworks that combine stable similarity metrics, fine-grained alignment, and cross-modal semantic supervision to enable scalable and robust generalization of Wi-Fi sensing.
@@ II-C3 Metric Learning for Zero/Few-shot Learning
Figure 11: Siamese Neural Networks. (1) The network takes a pair of inputs (e.g., two CSI segments) and feeds them into two identical sub-networks. Both branches share the same parameters (weights), mapping the inputs into a common low-dimensional feature space. (2) The model computes a similarity metric (usually Euclidean distance or Cosine similarity) between the two generated feature vectors. (3) Based on the distance, the system determines whether the two inputs belong to the same class (small distance) or different classes (large distance), often using a contrastive loss function.
Siamese neural networks [108] offer a classical yet effective architecture for few-shot learning and domain generalization. Unlike contrastive learning, which refers to a training paradigm defined by similarity-based objectives (e.g., contrastive or triplet loss), Siamese networks represent a weight-sharing architecture that is commonly used to implement such metric learning objectives. In this framework, as shown in Fig. 11, a pair of samples is fed into two identical networks that share weights. Each network extracts feature embeddings from its input, and the resulting embeddings are concatenated and passed through a binary classifier. The classifier is trained to predict whether the two inputs belong to the same class: if they do (e.g., two samples of the same activity), the target label is 1; otherwise, it is 0. During inference, a support set is constructed from a few labeled samples per class in the target domain. To classify a query sample, it is paired with each support sample and processed by the Siamese network. The network outputs a confidence score indicating the likelihood of the two samples belonging to the same class. The class label of the support sample with the highest confidence score is then assigned to the query sample.
This pipeline has been widely adopted in Wi-Fi sensing tasks for few-shot cross-domain generalization. For instance, 3D-ID [172] utilizes this pipeline for few-shot user identification in unseen environments. MaP-SGAN [128] applies it for gait recognition by enhancing the diversity of training samples via generative models. CrossFi [308] extends the Siamese framework to handle both cross-domain and new-class scenarios, enabling gesture recognition and user identification in previously unseen domains.
@@
Figure 12: Relation Network. (1) Both the support set and the query sample are passed through an embedding backbone to extract their respective feature maps. (2)The query feature map is paired and concatenated with each feature map in the support set, creating combined feature representations. (3) These concatenated pairs are fed into a Relation Module (usually a non-linear CNN or MLP), which acts as a learnable metric to calculate a relation score (between 0 and 1) for each pair. (4) The query is assigned to the class with the highest relation score, representing the strongest similarity determined by the deep neural network rather than a fixed distance metric.
@@
Figure 13: Matching Network. (1) Both the small support set and the query sample are mapped into a shared low-dimensional feature space using a neural network. (2) The model computes the cosine similarity (or distance metric) between the query embedding and each embedding in the support set. (3) These similarity scores are processed through a softmax function to generate attention weights, indicating which support examples most closely resemble the query. (4) The final label is predicted by a weighted sum of the support set labels, allowing the model to recognize new classes without any parameter updates.
@@
Figure 14: Prototypical Network. (1) For each class in the support set, the model computes a prototype (mean vector) by averaging the feature embeddings of all samples belonging to that class. (2) The query sample is mapped into the same feature space using the shared embedding backbone. (3) The model calculates the Euclidean distance between the query embedding and each class prototype. (4) The query is classified based on a softmax distribution over the negative distances; it is assigned to the class whose prototype is nearest in the embedding space.
@@
Summary and Design Insights for Metric Learning
Metric learning approaches share a common objective of structuring the embedding space to enable instance-level comparison rather than fixed classifier-based prediction, making them naturally suited for zero-shot and few-shot Wi-Fi sensing scenarios. Unlike domain alignment methods that primarily focus on reducing distribution discrepancy, metric learning emphasizes relational structure within the feature space, allowing the model to generalize to new domains or classes through support-set-based inference.
Among different paradigms, triplet-loss and contrastive-learning-based methods directly optimize pairwise relationships and often produce highly discriminative embeddings. They are flexible and compatible with various backbone architectures, but their performance depends on effective sample mining strategies and sufficient diversity within mini-batches. Siamese networks provide a straightforward architectural implementation of similarity learning, offering interpretability and ease of deployment, though they may suffer from scalability issues when the number of support comparisons grows.
Relation networks extend metric learning by introducing a learnable non-linear similarity function, enabling more expressive relational reasoning beyond fixed distance metrics. This improves adaptability to complex intra-class variations but increases computational overhead and model complexity. Matching networks and prototypical networks, in contrast, adopt episodic training to mimic test-time inference. Prototypical networks are particularly attractive due to their simplicity, computational efficiency, and strong performance under limited support samples, while matching networks provide a more flexible attention-based weighting mechanism at the cost of increased inference complexity.
From a design perspective, the choice of metric learning strategy should consider (i) the expected number of support samples per class, (ii) whether novel classes or only novel domains are encountered, (iii) computational constraints at inference time, and (iv) the degree of intra-class variability. Prototypical networks are well-suited for practical few-shot deployment with limited support data and strict efficiency requirements. Contrastive and triplet-based methods offer stronger embedding discrimination when large-scale pretraining is feasible. Relation-based models are preferable when complex relational reasoning is required, such as simultaneous cross-domain and new-class generalization.
Figure 15: Model-Agnostic Meta-Learning (MAML) [53] includes a meta-training stage and a meta-testing stage. The meta-training stage is designed to find a good initialization of model parameters such that the model can quickly adapt to new tasks with only a few gradient steps and a small amount of labeled data in the meta-testing stage.